OpenStudy (anonymous):

With the restriction x ≠ 2, which of the following is the simplified form of (3x-4/x-2) - (x-8/x-2) ? A)1 B)4(x+3)/x-2 C)x-2 D)2(x+6)/x-2

OpenStudy (dumbcow):

[(3x-4) - (x-8)]/x-2 = [2x +4]/x-2 = 2(x+2)/(x-2)

OpenStudy (dumbcow):

not one of the answers ???

OpenStudy (anonymous):

also \[3x-4/x-2 \neq (3x-4)/(x-2) \]